OpenStudy (anonymous):

I have filled a bag with differently colored ping-pong balls. There are twenty balls in the bag. Twelve are white, six are orange and two are pink. When I draw a ball from the bag, I remove it. What is the probability that my first draw is a white ball and my second is not?

OpenStudy (anonymous):

Ok...so: When you draw first ball, there are 12 white balls and a total of 20 balls So probability of drawing white ball = 12/20 = 3/5 When you draw second ball, there are 8 non-white balls and a total of 19 balls So probability of drawing non-white ball = 8/19 Probability of drawing white ball, then non-white ball = 3/5 * 8/19 = 24/95 = 0.252631579 = 25.26%

OpenStudy (anonymous):

thanks I had no clue how to even start that question